OpenStudy (anonymous):

Can someone pleeease help me simplify this expression.?? -6+i over -5+i

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (zzr0ck3r):

multiply top and bottom by (-5-i)

OpenStudy (anonymous):

I'm not sure what I get when I'm done with that...

OpenStudy (zzr0ck3r):

(-6+i)/(-5+i) = (-6+i)(-5-i)/((-5+i)(-5-i) = (30+i+1)/(25+1) = (31+i)/(26)

OpenStudy (anonymous):

Thank yoou soooo much!! Can you help me with another one?